OpenStudy (anonymous):

use implicit differentiation to find dy/dx square root of x +square root of y =1

OpenStudy (bahrom7893):

did i not just answer these?

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{y}=1\] \[\frac{1}{2\sqrt{x}}+\frac{1}{2\sqrt{y}}y'=0\] solve for y'

OpenStudy (anonymous):

\[y'=-\frac{\sqrt{y}}{\sqrt{x}}\]